OpenStudy (anonymous):

10/x -14=16, solve for x

OpenStudy (anonymous):

so first what you want to do is add 14 to both sides so 10/x = 30 now multiply by x on both sides 10 = 30x now divide by 30 on both sides 10/30 = x 1/3 = x

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

np